# Steinmetz Neural Networks for Complex-Valued Data

### v2

@@ Abstract
In this work, we introduce a new approach to processing complex-valued data using DNNs consisting of parallel real-valued subnetworks with coupled outputs. Our proposed class of architectures, referred to as Steinmetz Neural Networks, leverages multi-view learning to construct more interpretable representations within the latent space. Moreover, we present the Analytic Neural Network, which incorporates a consistency penalty that encourages analytic signal representations in the latent space of the Steinmetz neural network. This penalty enforces a deterministic and orthogonal relationship between the real and imaginary components. Utilizing an information-theoretic construction, we demonstrate that the generalization error upper bound posited by the analytic neural network is lower than that of the general class of Steinmetz neural networks. Our numerical experiments depict the improved performance and robustness to additive noise, afforded by these networks on benchmark datasets and synthetic examples.
@@ 1 Introduction
In recent years, the advancement of neural networks has spurred a wealth of research into specialized models designed for processing complex-valued data. Complex-valued neural networks (CVNNs) are a pivotal area of focus due to their intrinsic capability to leverage both the magnitude information and the phase information embedded in complex-valued signals, offering a distinct advantage over their real-valued counterparts (RVNNs) (Hirose, 2012; Guberman, 2016; Trabelsi et al., 2018). This is crucial across a spectrum of applications including telecommunications, medical imaging, and radar and sonar signal processing (Virtue et al., 2017; Gao et al., 2019; Smith, 2023). However, CVNNs are often encumbered by higher computational costs and more complex training dynamics (Bassey et al., 2021; Lee et al., 2022; Wu et al., 2023). These difficulties arise from the necessity to manage and optimize parameters within the complex domain, which can lead to instability in gradient descent methods and challenges in network convergence. Additionally, the search for effective and efficient complex-valued activation functions is a challenge (Scardapane et al., 2020; Lee et al., 2022).
While much of the discussion comparing CVNNs and RVNNs has been focused on the theoretical aspects of CVNNs, the development of improved RVNN architectures for complex-valued data processing remains an open problem. We propose addressing this problem from a feature learning perspective, leveraging multi-view representation fusion (Sun, 2013; Lahat et al., 2015; Zhao et al., 2017). By considering the independent and joint information of real and imaginary parts in successive processing steps, we aim to better capture the task-relevant information in complex-valued data.
In response to these considerations, this paper introduces the Steinmetz Neural Network architecture, a real-valued neural network that leverages multi-view learning to improve the processing of complex-valued data for predictive tasks. This architecture aims to mitigate the challenges of training CVNNs while forming more interpretable latent space representations, and comprises separate subnetworks that independently filter the irrelevant information present within the real and imaginary components, followed by a joint processing step. We note that the task-relevant interactions between components are not lost during the separate processing step, as they are handled in joint processing.
@@
The organization of this paper is as follows. In Section 2, we review complex and analytic signal representations, and survey the related work from CVNN literature. In Section 3, we present the Steinmetz neural network architecture and discuss its theoretical foundations. In Section 4, we summarize the consistency constraint and provide generalization error bounds for the Steinmetz network. In Section 5, we introduce the analytic neural network and the Hilbert transform consistency penalty. In Section 6, we present empirical results on benchmark datasets for complex-valued multi-class classification and regression, and provide synthetic experiments. In Section 7, we summarize our work.
@@
    We propose a consistency constraint on the latent space of the Steinmetz neural network to obtain a smaller upper bound on the generalization error, and derive these generalization bounds.
@@ 2 Preliminaries
To motivate our framework, we first formally define complex and analytic signal representations, and review existing real-valued neural networks and complex-valued neural networks for complex signal processing. Let $\smash{\mathcal{U}=\{0,1,\ldots,N-1\}}$ and $\smash{\mathcal{V}=\{0,1,\ldots,M-1\}}$, where $N\in\mathbb{N}$ is the signal period and $M\in\mathbb{N}$ denotes the size of the training dataset, $s$. To characterize the uncertainty of these signal representations, we denote $X\in\mathbb{C}^{dN}$, $X_{R},X_{I}\in\mathbb{R}^{dN}$ as the features, $Y\in\mathbb{C}^{k}$, $Y_{R},Y_{I}\in\mathbb{R}^{k}$ as the labels, and $Z\in\mathbb{C}^{lN}$, $Z_{R},Z_{I}\in\mathbb{R}^{lN}$ as the latent variables with $R,I$ denoting the respective real and imaginary parts, where:
@@ 2.1.1 Analytic Signal Representation
Above, $\mathcal{H}\{X_{R}^{m}\}$ denotes the discrete Hilbert transform (DHT) of the random variable $X_{R}^{m}$, wherein $\mathcal{U}^{\prime}=\{0,2,4,\ldots,dN\}$ for odd $n$, and $\mathcal{U}^{\prime}=\{1,3,5,\ldots,dN\}$ for even $n$. The DHT introduces a phase shift of $-90^{\circ}$ to all positive frequency components of $X_{R}^{m}$ and $+90^{\circ}$ to all negative frequency components of $X_{R}^{m}$, establishing orthogonality between the real and imaginary parts of $X^{m}$. We note that ${}\mathcal{H}\{\cdot\}$ is bijective since it is invertible, wherein $-\mathcal{H}\{\mathcal{H}\{X_{R}^{m}\}\}=X_{R}^{m}$, $\forall m\in\mathcal{V}$.
@@ 3 Steinmetz Neural Networks
Reflecting on the inherent challenges and benefits of both RVNNs and CVNNs, we target a framework that leverages the simplicity in training offered by RVNNs while offering improved generalization in the processing of complex signals. Within this context, multi-view representation fusion emerges as a potential framework, proposing that different perspectives — or ‘views’ — of data can provide complementary information, thereby enhancing learning and generalization (Sun, 2013; Xu et al., 2013; Lahat et al., 2015; Zhao et al., 2017; Yan et al., 2021). Deriving from this principle, we introduce the Steinmetz Neural Network architecture, which is designed to process the real ($X_{R}^{m}$) and the imaginary ($X_{I}^{m}$) parts of complex signal representations as separate views before joint processing. We formalize how this architecture leverages the complementarity principle of multi-view learning in Section 3.1.
@@ 3.1 Theoretical Foundations
In the context of neural networks’ information-theoretic foundations, Tishby and Zaslavsky (2015) used the Data Processing Inequality to formalize the architecture illustrated in Figure 1. This architecture aligns with the classical RVNN architecture from the CVNN literature (Trabelsi et al., 2018). Here, $\xi(\cdot)$ denotes the true function, $h(\psi(\cdot))$ denotes the neural network, and $\smash{\hat{Y}}^{m}=h(\psi(X^{m}))$ are the predictions — accordingly, we have that $I(Y^{m};X^{m})\geq I(Y^{m};Z^{m})\geq I(Y^{m};\smash{\hat{Y}}^{m})$. Regarding practical implementation, the input space, $\smash{[[X_{R}^{m}]^{T},[X_{I}^{m}]^{T}]^{T}\in\mathbb{R}^{2dN}}$, is jointly processed by $\psi^{*}(\cdot)$, using individual channels for $X_{R}^{m}$ and $X_{I}^{m}$ to form the latent space, $\smash{[[Z_{R}^{m}]^{T},[Z_{I}^{m}]^{T}]^{T}\in\mathbb{R}^{2lN}}$, which is then jointly processed by $h^{*}(\cdot)$, using individual channels for $Z_{R}^{m}$ and $Z_{I}^{m}$, to obtain the output space, $\smash{\hat{Y}^{m}\in\mathbb{C}^{k}}$.
@@
Building upon this architecture, the Steinmetz neural network postulates the Markov chain depicted in Figure 2, where the random variables, $Z_{I}^{m}=f(X_{I}^{m})$ and $Z_{R}^{m}=g(Z_{R}^{m})$, are the respective outputs of the parallel subnetworks, $f(\cdot)$ and $g(\cdot)$, where $Z^{m}=Z_{R}^{m}+iZ_{I}^{m}$ denotes the latent representation, and $\smash{\hat{Y}}^{m}=h(Z^{m})$ denotes the predictions yielded by the shared network, $h(\cdot)$  [^1] For practical implementation, we mean center $Z_{R}^{m}$ and $Z_{I}^{m}$ as the final step before concatenation.. Per the construction in Figure 2, the Steinmetz neural network is given by:
@@
For the classical RVNN, we jointly process $(X_{R}^{m},X_{I}^{m})$ and aim to form $(Z_{R}^{m},Z_{I}^{m})=\psi^{*}(X_{R}^{m},X_{I}^{m})$, filtering out $(\Lambda_{R}^{m},\Lambda_{I}^{m})$, where $(Z_{R}^{m},Z_{I}^{m})$, is a sufficient statistic of $(X_{R}^{m},X_{I}^{m})$ with respect to $Y^{m}$. We then jointly process $(Z_{R}^{m},Z_{I}^{m})$ and aim to form the predictions, $\smash{\hat{Y}}^{m}=h^{*}(Z_{R}^{m},Z_{I}^{m})$, filtering out $\Gamma^{m}$. We refer to this approach as joint-only processing. For the Steinmetz neural network, we separately process $X_{R}^{m}$ and $X_{I}^{m}$, and aim to form $Z_{R}^{m}=g(X_{R}^{m})$ and $Z_{I}^{m}=f(X_{I}^{m})$, filtering out $\Lambda_{R}^{m}$ and $\Lambda_{I}^{m}$, where $(Z_{R}^{m},Z_{I}^{m})$, is a sufficient statistic of $(X_{R}^{m},X_{I}^{m})$ with respect to $Y^{m}$. Paralleling the joint-only processing case, we jointly process this latent space, aiming to form $\smash{\hat{Y}}^{m}=h^{*}(Z_{R}^{m},Z_{I}^{m})$. We refer to this approach as separate-then-joint processing.
@@
We note that $\mathbf{K}_{Z_{R}^{m},Z_{I}^{m}}=\mathbf{0}_{dN\times dN}$ in the separate-then-joint processing approach, since $f(\cdot)$ and $g(\cdot)$ do not consider interactions between $Z_{R}^{m}$ and $Z_{I}^{m}$. This is summarized in Section B.1 of the Appendix. To measure the magnitude of the data interactions across both approaches, we consider the $L_{p,q}$ norm, with $p,q\geq 1$, of $\mathbf{\Sigma_{J}}$ and $\mathbf{\Sigma_{S}}$. As $\mathbf{\Sigma_{J}}$ includes the aforementioned cross-covariance matrix of $Z_{R}^{m}$ and $Z_{I}^{m}$, it follows that $\|\mathbf{\Sigma_{J}}\|_{p,q}\geq\|\mathbf{\Sigma_{S}}\|_{p,q}$, as shown in Corollary 3.1.
@@
This higher norm indicates reduced interpretability, as the presence of cross-covariance terms implies that joint-only processing must not only handle $Z_{R}^{m}$ and $Z_{I}^{m}$ individually, but also their interactions, which can complicate the representation process. As such, $\mathbf{\Sigma_{S}}$ has a smaller $L_{p,q}$ norm and associated representational complexity, enabling the Steinmetz network to separately extract $Z_{R}^{m}$ and $Z_{I}^{m}$ without the added burden of accounting for interactions. This Steinmetz architecture can also be leveraged to obtain a smaller upper bound on the generalization error, as per Section 4.
@@ 4 Consistency Constraint
Per Eq (19), the optimal latent representation, $Z^{m}$, best captures relevant information from $X^{m}$ about $Y^{m}$ while also considering the influence of $S$ on the encoder parameters, $\theta_{\psi_{S}}$. We now consider the upper bound on the generalization error over the training dataset, $\Delta s$, which is an adapted version of the bound originally proposed by (Kawaguchi et al., 2023).
@@
The upper bound from Eq. (4.1) captures the tradeoff between how well the latent space encapsulates information about the labels, and how much the encoder overfits the training distribution, wherein smaller values of $[I(X^{m};Z^{m}|Y^{m})+I(S;\theta_{\psi_{S}})]$ yield a smaller upper bound on the generalization error. Accordingly, we pose the following inquiry: is it possible to leverage the Steinmetz neural network architecture to obtain a smaller upper bound on the generalization error? To this end, we establish the existence of a lower bound, $\mathcal{D}(Z^{m})$, on $[I(X^{m};Z^{m}|Y^{m})+I(S;\theta_{\psi_{S}})]$ that is achievable using a constraint on the latent space of the Steinmetz neural network. We formalize this in Corollary 4.2.
@@
Achieving a smaller upper bound on the generalization error is indicative of a network’s potential for improved accuracy in predictions on unseen data. This relationship is deeply rooted in the notions of statistical learning theory, and more formally, in Structural Risk Minimization (SRM) and VC theory (Vapnik and Chervonenkis, 1971; Vapnik, 1999). Consequently, should there exist a consistency constraint on the latent space yielding a smaller upper bound on $\Delta s$, we would expect it to improve the Steinmetz neural network’s capacity to generalize (Vapnik, 2013).
@@
Theorem 4.3 informs us that $\mathcal{D}(Z^{m})$ is achievable when we enforce $Z_{I}^{m}=\phi(Z_{R}^{m})$, where $\phi(\cdot)$ is a deterministic, bijective function. We note that as the Steinmetz neural network is trained to minimize the average loss on the training dataset (via empirical risk minimization), we expect $Z^{m}$ to become more informative about the labels, whereby $I(Y^{m};Z^{m})$ increases. We now further extend this result to Theorem 4.1, via which we obtain a smaller upper bound on the generalization error.

### v3

@@ Abstract
We introduce a new approach to processing complex-valued data using DNNs consisting of parallel real-valued subnetworks with coupled outputs. Our proposed class of architectures, referred to as Steinmetz Neural Networks, incorporates multi-view learning to construct more interpretable representations in the latent space. Moreover, we present the Analytic Neural Network, which incorporates a consistency penalty that encourages analytic signal representations in the latent space of the Steinmetz neural network. This penalty enforces a deterministic and orthogonal relationship between the real and imaginary components. Using an information-theoretic construction, we demonstrate that the generalization gap upper bound posited by the analytic neural network is lower than that of the general class of Steinmetz neural networks. Our numerical experiments depict the improved performance and robustness to additive noise, afforded by our proposed networks on benchmark datasets and synthetic examples.
@@ 1 Introduction
In recent years, the advancement of neural networks has spurred a wealth of research into specialized models designed for processing complex-valued data. Complex-valued neural networks (CVNNs) are a pivotal area of focus due to their intrinsic capability to leverage both the magnitude information and the phase information embedded in complex-valued signals, offering a distinct advantage over their real-valued counterparts (RVNNs) (Hirose, 2012; Guberman, 2016; Trabelsi et al., 2018). This is crucial across a spectrum of applications including telecommunications, medical imaging, and radar and sonar signal processing (Virtue et al., 2017; Gao et al., 2019; Smith, 2023). However, CVNNs are often encumbered by higher computational costs and more complex training dynamics (Bassey et al., 2021; Lee et al., 2022; Wu et al., 2023). These difficulties arise from the necessity to manage and optimize parameters in the complex domain, which can lead to instability in gradient descent methods and challenges in network convergence. Additionally, the search for effective and efficient complex-valued activation functions is a challenge (Scardapane et al., 2020; Lee et al., 2022).
While much of the discussion comparing CVNNs and RVNNs has been focused on the theoretical aspects of CVNNs, the development of improved RVNN architectures for the processing of complex-valued data remains an open problem. We propose to address this problem through a feature learning perspective, leveraging multi-view representation fusion (Sun, 2013; Lahat et al., 2015; Zhao et al., 2017). By considering the independent and joint information of real and imaginary parts in successive processing steps, we aim to better capture the task-relevant information in complex-valued data.
In response to these considerations, this paper introduces the Steinmetz Neural Network architecture, a real-valued neural network that incorporates multi-view learning to improve the processing of complex-valued data for predictive tasks. This architecture aims to mitigate the challenges of training CVNNs while forming more interpretable latent space representations, and comprises separate subnetworks that independently filter the irrelevant information present within the real and imaginary components, followed by a joint processing step. We note that the task-relevant interactions between components are not lost during the separate processing step, as they are handled in joint processing.
@@
The organization of this paper is as follows. In Section 2, we review complex and analytic signal representations, and survey the related work from CVNN literature. In Section 3, we present the Steinmetz neural network architecture and discuss its theoretical foundations. In Section 4, we summarize the consistency constraint and provide generalization gap bounds for the Steinmetz network. In Section 5, we introduce the analytic neural network and the Hilbert transform consistency penalty. In Section 6, we present empirical results on benchmark datasets for complex-valued multi-class classification and regression, and provide synthetic experiments. In Section 7, we summarize our work.
@@
    We propose a consistency constraint on the latent space of the Steinmetz neural network to obtain a smaller upper bound on the generalization gap, and derive these generalization bounds.
@@ 2 Preliminaries
To motivate our framework, we begin by formally defining complex and analytic signal representations, and review existing real-valued neural networks and complex-valued neural networks for complex signal processing. Let $\smash{\mathcal{U}=\{0,1,\ldots,N-1\}}$ and $\smash{\mathcal{V}=\{0,1,\ldots,M-1\}}$, where $N\in\mathbb{N}$ is the signal period and $M\in\mathbb{N}$ denotes the size of the training dataset, $s$. To characterize the uncertainty of these signal representations, we denote $X\in\mathbb{C}^{dN}$, $X_{R},X_{I}\in\mathbb{R}^{dN}$ as the features, $Y\in\mathbb{C}^{k}$, $Y_{R},Y_{I}\in\mathbb{R}^{k}$ as the labels, and $Z\in\mathbb{C}^{lN}$, $Z_{R},Z_{I}\in\mathbb{R}^{lN}$ as the latent variables with $R,I$ denoting the respective real and imaginary parts, where:
@@ 2.1.1 Analytic Signal Representation
$\mathcal{H}\{X_{R}^{m}\}$ is the discrete Hilbert transform (DHT) of $X_{R}^{m}$, where $\mathcal{U}^{\prime}=\{\,u\in\{0,1,\dots,dN-1\},u\not\equiv n\pmod{2}\,\}$, which selects even indices when $n$ is odd and odd indices when $n$ is even. The DHT introduces a phase shift of $-90^{\circ}$ to all the positive frequency components of $X_{R}^{m}$ and $+90^{\circ}$ to all the negative frequency components of $X_{R}^{m}$, establishing orthogonality between $X_{R}^{m}$ and $X_{I}^{m}$. We observe that ${}\mathcal{H}\{\cdot\}$ is bijective since it is invertible, wherein $-\mathcal{H}\{\mathcal{H}\{X_{R}^{m}\}\}=X_{R}^{m}$, $\forall m\in\mathcal{V}$.
@@ 3 Steinmetz Neural Networks
Reflecting on the inherent challenges and benefits of both RVNNs and CVNNs, we target a framework that leverages the simplicity in training offered by RVNNs while offering improved generalization in the processing of complex signals. Within this context, multi-view representation fusion emerges as a potential framework, proposing that different perspectives — or ‘views’ — of data can provide complementary information, thereby enhancing learning and generalization (Sun, 2013; Xu et al., 2013; Lahat et al., 2015; Zhao et al., 2017; Yan et al., 2021). Informed by this principle, we introduce the Steinmetz Neural Network architecture, which is designed to process the real ($X_{R}^{m}$) and the imaginary ($X_{I}^{m}$) parts of complex signal representations as separate views before joint processing. We formalize how this architecture leverages the complementarity principle of multi-view learning in Section 3.1.
@@ 3.1 Theoretical Foundations
In the context of neural networks’ information-theoretic foundations, (Tishby and Zaslavsky, 2015) leveraged the Data Processing Inequality to characterize the architecture illustrated in Figure 1. This architecture aligns with the classical RVNN architecture from the literature on CVNNs (Trabelsi et al., 2018). Here, $\xi(\cdot)$ denotes the true function, $h(\psi(\cdot))$ describes the neural network, and $\smash{\hat{Y}}^{m}=h(\psi(X^{m}))$ denotes the predictions. Accordingly, we have that $I(Y^{m};X^{m})\geq I(Y^{m};Z^{m})\geq I(Y^{m};\smash{\hat{Y}}^{m})$. Regarding practical implementation, the input space, $\smash{[[X_{R}^{m}]^{T},[X_{I}^{m}]^{T}]^{T}\in\mathbb{R}^{2dN}}$, is jointly processed by $\psi^{*}(\cdot)$, using individual channels for $X_{R}^{m}$ and $X_{I}^{m}$ to form the latent space, $\smash{[[Z_{R}^{m}]^{T},[Z_{I}^{m}]^{T}]^{T}\in\mathbb{R}^{2lN}}$, which then gets jointly processed by $h^{*}(\cdot)$, using individual channels for $Z_{R}^{m}$ and $Z_{I}^{m}$, to obtain the output space, $\smash{\hat{Y}^{m}\in\mathbb{C}^{k}}$.
@@
Building upon this architecture, the Steinmetz neural network postulates the Markov chain depicted in Figure 2, where the random variables, $Z_{I}^{m}=f(X_{I}^{m})$ and $Z_{R}^{m}=g(Z_{R}^{m})$, are the respective outputs of the parallel subnetworks, $f(\cdot)$ and $g(\cdot)$, where $Z^{m}=Z_{R}^{m}+iZ_{I}^{m}$ denotes the latent representation, and $\smash{\hat{Y}}^{m}=h(Z^{m})$ denotes the predictions yielded by the shared network, $h(\cdot)$  [^1] footnotetext: For practical implementation, we mean center $Z_{R}^{m}$ and $Z_{I}^{m}$ as the final step before concatenation.. Per the construction in Figure 2, the Steinmetz neural network is given by:
@@
For the classical RVNN, we jointly process $(X_{R}^{m},X_{I}^{m})$ and aim to form $(Z_{R}^{m},Z_{I}^{m})=\psi^{*}(X_{R}^{m},X_{I}^{m})$, filtering out $(\Lambda_{R}^{m},\Lambda_{I}^{m})$, where $(Z_{R}^{m},Z_{I}^{m})$, is a sufficient statistic of $(X_{R}^{m},X_{I}^{m})$ with respect to $Y^{m}$. We then jointly process $(Z_{R}^{m},Z_{I}^{m})$ and attempt to form the predictions, $\smash{\hat{Y}}^{m}=h^{*}(Z_{R}^{m},Z_{I}^{m})$, filtering out $\Gamma^{m}$. We refer to this approach as joint-only processing. For the Steinmetz neural network, we separately process $X_{R}^{m}$ and $X_{I}^{m}$, and attempt to form $Z_{R}^{m}=g(X_{R}^{m})$ and $Z_{I}^{m}=f(X_{I}^{m})$, filtering out $\Lambda_{R}^{m}$ and $\Lambda_{I}^{m}$, where $(Z_{R}^{m},Z_{I}^{m})$, is a sufficient statistic of $(X_{R}^{m},X_{I}^{m})$ with respect to $Y^{m}$. Paralleling the joint-only processing case, we jointly process this latent space, aiming to form $\smash{\hat{Y}}^{m}=h^{*}(Z_{R}^{m},Z_{I}^{m})$. We refer to this approach as separate-then-joint processing.
@@
We note that $\mathbf{K}_{Z_{R}^{m},Z_{I}^{m}}=\mathbf{0}_{dN\times dN}$ in the separate-then-joint processing approach, since $f(\cdot)$ and $g(\cdot)$ do not consider the interactions between $Z_{R}^{m}$ and $Z_{I}^{m}$. This is characterized in Section B.1 of the Appendix. To measure the magnitude of the data interactions across both approaches, we consider the $L_{p,q}$ norm, with $p,q\geq 1$, of $\mathbf{\Sigma_{J}}$ and $\mathbf{\Sigma_{S}}$. As $\mathbf{\Sigma_{J}}$ includes the aforementioned cross-covariance matrix of $Z_{R}^{m}$ and $Z_{I}^{m}$, it follows that $\|\mathbf{\Sigma_{J}}\|_{p,q}\geq\|\mathbf{\Sigma_{S}}\|_{p,q}$, as shown in Corollary 3.1.
@@
This greater norm indicates reduced interpretability, as the presence of cross-covariance terms implies that joint-only processing must not only handle $Z_{R}^{m}$ and $Z_{I}^{m}$ individually, but also their interactions, which can complicate the representation process. As such, $\mathbf{\Sigma_{S}}$ has a smaller $L_{p,q}$ norm and associated representational complexity, enabling the Steinmetz network to separately extract $Z_{R}^{m}$ and $Z_{I}^{m}$ without the added burden of accounting for interactions. This Steinmetz architecture can also be leveraged to obtain a smaller upper bound on the generalization gap, as per Section 4.
@@ 4 Consistency Constraint
Per Eq (19), the optimal latent representation, $Z^{m}$, best captures relevant information from $X^{m}$ about $Y^{m}$ while also considering the influence of $S$ on the encoder parameters, $\theta_{\psi_{S}}$. We now consider the upper bound on the generalization gap over the training dataset, $\Delta s$, which is an adapted version of the bound originally proposed in (Kawaguchi et al., 2023).
@@
The upper bound from Eq. (4.1) captures the tradeoff between how well the latent space encapsulates information about the labels, and how much the encoder overfits the training distribution, wherein smaller values of $[I(X^{m};Z^{m}|Y^{m})+I(S;\theta_{\psi_{S}})]$ yield a smaller upper bound on the generalization gap. Accordingly, we pose the following inquiry: is it possible to leverage the Steinmetz neural network architecture to obtain a smaller upper bound on the generalization gap? To this end, we establish the existence of a lower bound, $\mathcal{D}(Z^{m})$, on $[I(X^{m};Z^{m}|Y^{m})+I(S;\theta_{\psi_{S}})]$ that is achievable using a constraint on the latent space of the Steinmetz neural network. We formalize this in Corollary 4.2.
@@
Achieving a smaller upper bound on the generalization gap is indicative of a network’s potential for improved accuracy in making predictions on unseen data. This relationship is deeply rooted in the notions of statistical learning theory, and more formally, in Structural Risk Minimization (SRM) and VC theory (Vapnik and Chervonenkis, 1971; Vapnik, 1999). Consequently, should there exist a consistency constraint on the latent space yielding a smaller upper bound on $\Delta s$, we would expect it to improve the Steinmetz neural network’s capacity to generalize (Vapnik, 2013).
@@
Theorem 4.3 informs us that $\mathcal{D}(Z^{m})$ is achievable when we enforce $Z_{I}^{m}=\phi(Z_{R}^{m})$, where $\phi(\cdot)$ is a deterministic, bijective function. We note that as the Steinmetz neural network is trained to minimize the average loss on the training dataset (through empirical risk minimization), we expect $Z^{m}$ to become more informative about the labels, whereby $I(Y^{m};Z^{m})$ increases. We now further extend this result to Theorem 4.1, via which we obtain a smaller upper bound on the generalization gap.